Two cars leave philadelphia, one traveling north and the other traveling south. After 3 hours, they are 297 miles apart. If one car is traveling 5 miles per hour faster than the other, what is the speed of each?
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

Two cars leave philadelphia, one traveling north and the other traveling south. After 3 hours, they are 297 miles apart. If one car is traveling 5 miles per hour faster than the other, what is the speed of each?
-----------------------------------
Slower car DATA:
rate = x mph ; time = 3 hrs ; distance = 3x miles
-----------------
Faster car DATA:
rate = x+5 mph ; time = 3 hrs ; distance = 3x+15 miles
---------------------
Equation:
3x + 3x+15 = 297
6x + 15 = 297
2x + 5 = 99
2x = 94
x = 47 mph (slower car rate)
x+5 = 52 (faster car rate)
==============================
Cheers,
Stan H.
==========================

Answer by checkley77(12844)

D=RT
297=(X+X+5)3
297=(2X+5)3
297=6X+15
6X=297-15
6X=282
X=282/6
X=47 MPH. FOR THE SLOWER CAR.
47+5=52 MPH. FOR THE FASTER CAR.
PROOF:
297=(47+52)3
297=99*3
297=297